# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v1

@@ Abstract
Evolutionarily stable strategy (ESS) is the defining concept of evolutionary game theory. It has a fairly unanimously accepted definition for the case of symmetric games which are played in a homogeneous population where all individuals are in same role. However, in asymmetric games, which are played in a population with multiple subpopulations (each of which has individuals in one particular role), situation is not as clear. Various generalizations of ESS defined for such cases differ in how they correspond to fixed points of replicator equation which models evolutionary dynamics of frequencies of strategies in the population. Moreover, some of the definitions may even be equivalent, and hence, redundant in the scheme of things. Along with reporting some new results, this paper is partly indented as a contextual mini-review of some of the most important definitions of ESS in asymmetric games. We present the definitions coherently and scrutinize them closely while establishing equivalences—some of them hitherto unreported—between them wherever possible. Since it is desirable that a definition of ESS should correspond to asymptotically stable fixed points of replicator dynamics, we bring forward the connections between various definitions and their dynamical stabilities. Furthermore, we find the use of principle of relative entropy to gain information-theoretic insights into the concept of ESS in asymmetric games, thereby establishing a three-fold connection between game theory, dynamical system theory, and information theory in this context. We discuss our conclusions also in the backdrop of asymmetric hypermatrix games where more than two individuals interact simultaneously in the course of getting payoffs.
@@ I Introduction
There have been multiple attempts to extend the concept of ESS to asymmetric and bimatrix games. Taylor Taylor (1979) defined asymmetric ESS as the resilient strategy that fetches more total average payoff—summation of average payoffs of every subpopulation of individuals with distinct role—than what a mutant strategy would do. Around the same time, Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, in terms of requiring that average payoffs in every role be more; interestingly, unlike the case of symmetric games, the ESS can not be a mixed strategy in bimatrix games Pohley and Thomas (1979); Schuster et al. (1981). Consequently, alternative definitions of ESS in asymmetric and bimatrix games were proposed to allow for the mixed ESS concept. For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that average payoff in any one of the role should be more than what corresponding mutant could get. This concept is termed two-species ESS—henceforth abbreviated as 2ESS. For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of ESS, again with a view to allow for emergence of mixed ESS. The concept is termed Nash–Pareto pair.
@@
The push for getting mixed ESS in asymmetric and bimatrix games stems from the fact that the replicator equation Taylor and Jonker (1978); Cressman and Tao (2014)—the evolutionary dynamics in line with the Darwinian tenet of evolution—possesses asymptotically stable interior fixed points. It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points, in other words, the states of the population. Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability in which cases. Most importantly, we find that the connection between 2ESS and its dynamical stability within the paradigm of replicator equation is not as exhaustive and extensive as was claimed earlier in the literature Smith and Price (1973); Smith (1982). Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
@@
The outline of the paper is as follows: We commence with Sec. II by outlining the existing definitions of ESS in bimatrix games. Rather than following a chronological description, we adopt a logical and coherent narrative suited for the goal of this paper. We demonstrate the equivalences between some of the definitions. Next, in Sec. III, we highlight caveats and issues associated with these definitions in bimatrix games through illustrative examples while using the concept of strong stability Hofbauer and Sigmund (2003). We understand ESS further from the angle of dynamical stability using replicator equation. Sec. IV presents the implications of introducing intra-specific interaction in bimatrix games, which essentially makes them asymmetric games. In Sec. V, we discuss some information-theoretic connections of ESS. In the next section, Sec. VI, we generalize hitherto obtained results to the case of multiplayer setting, i.e., where more than two players interact simultaneously. Finally, we conclude in Sec. VII.
@@ II Equivalence between various definitions of ESS in Bimatrix Games
Consider a population consisting of two subpopulations denoted as X and Y. Let $\mathbf{x}=(x_{1},x_{2},\cdots,x_{m})$ represent a state of subpopulation X, where each $x_{i}$ represents the frequency of the $i$th phenotype in the subpopulation. Similarly, let $\mathbf{y}=(y_{1},y_{2},\cdots,y_{n})$ represent a state of subpopulation Y, where each $y_{j}$ represents the frequency of the $j$th type in the subpopulation. The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$, which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$. In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff matrix of type $i$ in Y against type $j$ in X. All such payoffs are collectively expressed through the $m\times n$ matrix ${\sf B}$ with elements $b_{ij}$’s and and $n\times m$ matrix ${\sf C}$ with elements $c_{ij}$’s.
Let $(\mathbf{x},\mathbf{y})$ denote any arbitrary state of the population. Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionary stable state (ESS). As is the idea behind ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ may be said to be an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant. Since practically mutations are rather rare, it makes sense for this criterion to hold good for infinitesimal mutant fraction in the context of infinitely large host population. Because the host population has more than one subpopulation, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
@@ II.1 Mutants in both subpopulations
Let us begin with following set of three definitions which we shall show to be equivalent:
  1. Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$, with $0<\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>\mathbf{x}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})~{}~{}~{}\text{and}~{}~{}~{}\mathbf{\hat{y}}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{y}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (1)
    Here, $\epsilon_{1}$ and $\epsilon_{2}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
@@
    This reminds one of what Hofbauer and Sigmund termed the “weak ESS” Hofbauer and Sigmund (1998); it also includes the definitions Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed as ESS for bimatrix games.
@@
Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations. Schuster et al. (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979). Accordingly, we write the following set of three equivalent definitions:
  1. Definition 2a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called an ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$, with $0<\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})+\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{x}\cdot{\sf B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})+\mathbf{y}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (6)
  2. Definition 2b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
@@
It is obvious that when the Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true. We would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b) and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
@@ II.2 Mutants in both or either subpopulations
Few remarks are in order. ESS’s defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al. (1981). Hence, this fact prompted a need to revise the definitions so as to accommodate scenarios involving mixed ESS. The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and deem it adequate for any one resident to beat the mutants. In definitions 1 and 2, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants. Taylor Taylor (1979) discussed this point, highlighting that if mutants only arrive in subpopulation Y, one should compare the fitness of residents of Y versus mutants of Y while interacting with residents of population X and vice versa. He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in Definition 2. In the light of these comments, we now present following set of equivalent definitions where one of the $\epsilon_{1}$ or $\epsilon_{2}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
  1. Definition 3a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$ with $0\leq\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0\leq\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$$(\epsilon_{1},\epsilon_{2})\neq(0,0)$,
    $\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>\mathbf{x}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{y}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (10)
  2. Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,

### v2

@@ Abstract
Evolutionarily stable strategy (ESS) is the defining concept of evolutionary game theory. It has a fairly unanimously accepted definition for the case of symmetric games which are played in a homogeneous population where all individuals are in same role. However, in asymmetric games, which are played in a population with multiple subpopulations (each of which has individuals in one particular role), situation is not as clear. Various generalizations of ESS defined for such cases differ in how they correspond to fixed points of replicator equation which models evolutionary dynamics of frequencies of strategies in the population. Moreover, some of the definitions may even be equivalent, and hence, redundant in the scheme of things. Along with reporting some new results, this paper is partially indented as a contextual mini-review of some of the most important definitions of ESS in asymmetric games. We present the definitions coherently and scrutinize them closely while establishing equivalences—some of them hitherto unreported—between them wherever possible. Since it is desirable that a definition of ESS should correspond to asymptotically stable fixed points of replicator dynamics, we bring forward the connections between various definitions and their dynamical stabilities. Furthermore, we find the use of principle of relative entropy to gain information-theoretic insights into the concept of ESS in asymmetric games, thereby establishing a three-fold connection between game theory, dynamical system theory, and information theory in this context. We discuss our conclusions also in the backdrop of asymmetric hypermatrix games where more than two individuals interact simultaneously in the course of getting payoffs.
@@ I Introduction
There have been multiple attempts to extend the concept of ESS to asymmetric and bimatrix games. Taylor Taylor (1979) defined asymmetric ESS as the resilient strategy that fetches more total average payoff—summation of average payoffs of every subpopulation of individuals with distinct role—than what a mutant strategy would do. Around the same time, Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, in terms of requiring that average payoffs in every role be more; interestingly, unlike the case of symmetric games, the ESS can not be a mixed strategy in bimatrix games Pohley and Thomas (1979); Schuster et al. (1981). Consequently, alternative definitions of ESS in asymmetric and bimatrix games were proposed to allow for the mixed ESS concept. For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that average payoff in any one of the role should be more than what corresponding mutant could get. This concept is termed two-species ESS—henceforth abbreviated as 2ESS. For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of 2ESS, again with a view to allow for emergence of mixed ESS. The concept is termed Nash–Pareto pair.
@@
The push for getting mixed ESS in asymmetric and bimatrix games stems from the fact that the replicator equation Taylor and Jonker (1978); Cressman and Tao (2014)—the evolutionary dynamics in line with the Darwinian tenet of evolution—possesses asymptotically stable interior fixed points. It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points, in other words, the states of the population. Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability in which cases. Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
@@
The outline of the paper is as follows: We commence with Sec. II by outlining the existing definitions of ESS in bimatrix games. Rather than following a chronological description, we adopt a logical and coherent narrative suited to the goal of this paper. We demonstrate the equivalences between some of the definitions. Next, in Sec. III, we highlight caveats and issues associated with these definitions in bimatrix games. We scrutinize the concept of ESS further from the angle of dynamical stability under replicator dynamics. Sec. IV presents the implications on bimatrix games due to intra-specific interactions which essentially turns them into general asymmetric games. In Sec. V, we discuss some interesting information-theoretic connections of ESS. In the next section, Sec. VI, we generalize hitherto obtained results for multiplayer hypermatrix games. Finally, we conclude in Sec. VII.
@@ II Equivalence between various definitions of ESS in bimatrix Games
Consider a population consisting of two subpopulations denoted as X and Y. Let $\mathbf{x}=(x_{1},x_{2},\cdots,x_{m})$ represent a state of subpopulation X, where each $x_{i}$ represents the frequency of the $i$th phenotype in the subpopulation. Similarly, let $\mathbf{y}=(y_{1},y_{2},\cdots,y_{n})$ represent a state of subpopulation Y, where each $y_{j}$ represents the frequency of the $j$th type in the subpopulation. The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$, which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$. In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff of type $i$ in Y against type $j$ in X. All such payoffs are collectively expressed through the $m\times n$ matrix ${\sf B}$ with elements $b_{ij}$’s and and $n\times m$ matrix ${\sf C}$ with elements $c_{ij}$’s.
Let $(\mathbf{x},\mathbf{y})$ denote any arbitrary state of the population. Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionarily stable state (ESS). As is the idea behind ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ may be said to be an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant. Since practically mutations are rather rare, it makes sense for this criterion to hold good for infinitesimal mutant fraction in the context of infinitely large host population. Because the host population has more than one subpopulation, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
@@ II.1 Mutants in both subpopulations
Let us begin with following set of three definitions (aptly termed and-definition) which can be shown to be equivalent:
  1. Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{and}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (1)
    Here, $\epsilon_{x}$ and $\epsilon_{y}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
@@
    This reminds one of what Hofbauer and Sigmund termed the “weak ESS” Hofbauer and Sigmund (1998); it also includes the definitions Selten Selten (1980) and Schuster et al. Schuster et al. (1981); Hofbauer and Sigmund (1988) proposed as ESS for bimatrix games.
@@
It is of use to recall that a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a Nash equilibrium Nash (1950, 1951); Hofbauer and Sigmund (1998) (NE) in bimatrix games if the following is true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$:
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}~{}~{}\text{and}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}.$ (6)
Mere inspection reveals that NE is directly implied by and-definition; and, of course, when NE is strict, ESS under and-definition is implied.
Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations Schuster et al. (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979). Accordingly, we write the following set of three equivalent definitions (aptly termed sum-definition):
  1. Definition 2a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called an ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (7)
  2. Definition 2b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
@@
It is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true. We would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b) and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
The equilibrium condition of Definition 2b (Eq. (8)) appears different from Eq. (6). So while strict NE obviously implies ESS under sum-definition, it is not obvious if the ESS implies NE always. To check this out, let’s write inequality (8) in the following form: $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$. Since this has to be true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, if $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}<0$ for some $\mathbf{x}$, then one can always find a $\mathbf{y}$ such that $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}<0$ which means both the terms individually has to satisfy $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}\geq 0$ and $(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$ for all $({\bf x},{\bf y})$ so that inequality (8) holds good. Concluding, an ESS under sum-definition implies NE.
We emphasize that the definitions in this section require simultaneous appearance of mutants in both the subpopulations, i.e., $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ as mentioned in Definitions 1a and 2a explicitly. In other words, states $(\mathbf{\hat{x}},\mathbf{y})$ and $(\mathbf{x},\mathbf{\hat{y}})$ are not mathematically allowed mutant states. This restriction is relaxed in the next definition as introduced in the next section.
@@ II.2 Mutants in both or either subpopulations
Few remarks are in order. ESS’s defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al. (1981). Hence, this fact prompted a need to revise the definitions so as to accommodate scenarios involving mixed ESS. The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and deem it adequate for any one resident to beat the mutants. In and- and sum-definitions, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants. Taylor Taylor (1979) discussed this point, highlighting that if mutants only arrive in subpopulation Y, one should compare the fitness of residents of Y versus mutants of Y while interacting with residents of population X and vice versa. He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in sum-definition. In the light of above remarks, we now present the following set of equivalent definitions (aptly termed or-definition) where one of $\epsilon_{x}$ or $\epsilon_{y}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
  1. Definition 3a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$ with $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$$(\epsilon_{x},\epsilon_{y})\neq(0,0)$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{or}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (11)
  2. Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
@@
    (It basically captures the fact that both the mutants cannot simultaneously take advantage of the resident; at least one of them gets penalized.)